OpenStudy (anonymous):

As the marketing designer for Burger King, you are given a 4 in by 6 in logo and have been told to create an advertisement with side lengths six times larger. What is the area of your advertisement (in square feet)? [12 inches = 1 ft] A. 6 ft2 B. 36 ft2 C. 72 ft2 D. 3 ft2

OpenStudy (anonymous):

c because 4*6=24 and 6*6=36 and 24*36=864 864/12=72

OpenStudy (anonymous):

its 6 feet

OpenStudy (anonymous):

my bad it is 72 sorry

OpenStudy (anonymous):

thats ok